OpenStudy (goformit100):

Two bus tickets from city A to B and three tickets from city A to C cost Rs. 77 but three tickets from city A to B and two tickets from city A to C cost Rs. 73. What are the fares for cities B and C from A ? A. Rs. 4, Rs. 23 B. Rs. 13, Rs. 17 C. Rs. 15, Rs. 14 D. Rs. 17, Rs. 13

OpenStudy (goformit100):

@Suchandrarao

OpenStudy (anonymous):

let ticket from A to B be x let ticket from A to C be y so 2x+3y=77 3x+2y=73 solve for x and y

OpenStudy (hoblos):

let B be the fares for city B from A let C be the fares for city C from A you can convert this to the following system 2B + 3C = 77 3B + 2C = 73 can you solve the system ?

OpenStudy (anonymous):

when u solve this the ans would be x=13, y=17